OpenStudy (anonymous):

A farm has a dimension of 85 m by 45 m. 3/5 of its area is planted with corn. the remaining area was planted with monggo. how wide is the area planted with monggo?

OpenStudy (anonymous):

A= 85 * 45 = 3825

OpenStudy (anonymous):

then the equivalent value of 3/5 is 0.6

OpenStudy (anonymous):

should i subtract it?

hartnn (hartnn):

oh, don't subtract it. since 3/5 of the are is planted, its 3/5 * total so AREA PLANTED = 3/5 * 3825 = ?

OpenStudy (anonymous):

2295?

hartnn (hartnn):

that is are planted with corn. now remaining area is planted with mango. so answer = 3825-2295= ?

OpenStudy (anonymous):

ahhhh now i get it.

OpenStudy (anonymous):

1530

hartnn (hartnn):

yup, thats it :)